OpenStudy (jgresham5):

What number is 32% of 393.75

OpenStudy (visarika23):

Let that no be x So 32% of x=393.75 (32/100)*x=393.75 x=(393.75*100)/32

OpenStudy (visarika23):

Did you get that??